OpenStudy (anonymous):

Hi. I think someone already asked this question, but I am really having a hard time with it. Functions f(x) and g(x) are shown below: f(x)=3x^2+12x+16 g(x)=2sin(2x-pi)+4 Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value.

OpenStudy (anonymous):

*Find critical poi.nts and use second derivative test F(x)=6x+12=0 -> x=-2 Since F(x)=6>0, minimum is at -2 g(x)=4cos(4x-pi)=0, -pi/4 -> x= pi/2, 3pi/2 etc... g(x)=16sin(4x-pi) >0, x= 5pi/8 etc... g(5pi/8)=2sin(pi/4)+4=4+ sqrt2 > F(-2)=4